# PKF: Probabilistic Data Association Kalman Filter for Multi-Object Tracking

### v1

@@ Abstract
In this paper, we derive a new Kalman filter with probabilistic data association between measurements and states. We formulate a variational inference problem to approximate the posterior density of the state conditioned on the measurement data. We view the unknown data association as a latent variable and apply Expectation Maximization (EM) to obtain a filter with update step in the same form as the Kalman filter but with expanded measurement vector of all potential associations. We show that the association probabilities can be computed as permanents of matrices with measurement likelihood entries. We also propose an ambiguity check that associates only a subset of ambiguous measurements and states probabilistically, thus reducing the association time and preventing low-probability measurements from harming the estimation accuracy. Experiments in simulation show that our filter achieves lower tracking errors than the well-established joint probabilistic data association filter (JPDAF), while running at comparable rate. We also demonstrate the effectiveness of our filter in multi-object tracking (MOT) on multiple real-world datasets, including MOT17, MOT20, and DanceTrack. We achieve better higher order tracking accuracy (HOTA) than previous Kalman-filter methods and remain real-time. Associating only bounding boxes without deep features or velocities, our method ranks top-10 on both MOT17 and MOT20 in terms of HOTA. Given offline detections, our algorithm tracks at 250+ fps on a single laptop CPU. Code is available at https://github.com/hwcao17/pkf.
1 Introduction
Figure 1: Example scene with high ambiguity in DanceTrack [40] with green box detections and orange box tracks. Applying our probabilistic data association Kalman filter on this sequence significantly increases the association quality according to IDF1 [38], association accuracy (AssA) [24], association recall (AssRe) [24], and the combined HOTA metric [24].
@@ 1 Introduction
In tasks where ambiguity exists between measurements and variables of interest, probabilistic data association can prevent catastrophic estimation failures. For example, in multi-object tracking (MOT), there can be a lot of occlusions causing high ambiguity in data association. An illustration of how probabilistic data association improves MOT is shown in Figure 1. Methods utilizing the Kalman filter [6, 47, 13] have achieved outstanding performance in MOT but have not considered the impact of probabilistic data association on the tracking process.
@@
We derive a new formulation of the Kalman filter with probabilistic data association. Previous works [3, 12] show that the Kalman filter can be derived using variational inference (VI) by maximizing the evidence lower bound (ELBO) of the input and measurement data likelihood. To deal with ambiguous associations, we approach the VI problem using Expectation-Maximization (EM) with the data association as the latent variable. In the E-step, we show that the association weights can be computed as permanents of matrices with measurement likelihood entries using accelerated permanent algorithms [32, 20]. We also show that the weight computation can be extended to cases with missing detections and clutter. In the M-step, optimizing the EM objective once leads to a Kalman filter with the usual prediction and update steps but with an extended measurement vector and noise covariance matrix capturing all possible data associations. Our formulation is related to the well-known joint probabilistic data association filter (JPDAF) [2]. However, while JPDAF first computes a posterior given each associated measurement, i.e., maximizes the ELBO of each measurement, and then computes the weighted average of the posteriors, our filter directly optimizes the overall ELBO weighted by the data association probabilities. We show that our algorithm achieves lower tracking errors than JPDAF, while running at comparable tracking rates.
We also apply our filter to MOT. We observe that, with accurate detection algorithms [36, 35], binary associations are fast and work well when the ambiguity is not severe and incorporating measurements with low association weights can harm the estimation accuracy. Therefore, we employ a hybrid data association procedure. We first do binary data association (e.g., via [23, 21]) and then perform an ambiguity check which outputs a set of ambiguous measurements and tracks. Then, probabilistic data association is only performed on the ambiguous set, thus reducing the computation time and preventing low-probability measurements from harming the accuracy. We test our algorithm on multiple real-world MOT datasets, including MOT17 [27], MOT20 [15], and DanceTrack [40]. Our method outperforms previous Kalman-filter methods [47, 13] while maintaining almost the same inference speed. We name our probabilistic data association Kalman filter the PKF with P emphasizing both the probabilistic nature of the data association and the matrix permanent computation of the association weights. Our contributions are summarized as follows.
@@
  • We show that association probabilities can be computed using matrix permanents and design a hybrid association procedure with an ambiguity check to reduce computation complexity and exclude low-probability associations.
  • We demonstrate the effectiveness of the PKF in comparison to the JPDAF and on multiple real-world MOT datasets. PKF ranks top-10 on both MOT17 and MOT20 by associating only bounding boxes and runs at 250+ fps on a single laptop CPU given offline detections.
2 Related work
Probabilistic data association
@@ 2 Related work
Multi-object tracking State-of-art MOT algorithms are dominated by the tracking-by-detection approach, which consists of two steps: object detection [36, 35] and data association. SORT [6] uses the Hungarian algorithm [23] to associate bounding-box object detections and updates the object tracks using the Kalman filter to achieve real-time online tracking. DeepSORT [42] extends SORT [6] by introducing deep visual features for data association. ByteTrack [47] makes an observation that the detectors can make imperfect predictions in complex scenes and proposes a second round of data association for low-confidence detections. OC-SORT [13] analyzes the limitations of SORT [13], namely sensitivity to state noise and temporal error magnification, and proposes an observation-centric re-update step and object momentum for data association.
@@
3 Problem formulation
Consider $N$ time-varying variables that need to be estimated, denoted as $\mathbf{x}_{t,j}\in\mathbb{R}^{n}$, $j=1,...,N$. We refer to their concatenation into a single vector at time $t$:
@@ 3 Problem formulation
where $\mathbf{u}_{t}$ is the known input and $\mathbf{w}_{t}$ is zero-mean Gaussian noise with covariance $W$.
At each time $t$, a subset of the variables $\mathbf{x}_{t,1},\ldots,\mathbf{x}_{t,N}$ are observed. Let $M_{t}\in\{0,1,\ldots,N\}$ be the number of measurements at time $t$ and let $\mathbf{z}_{t,k}\in\mathbb{R}^{m}$ be the $k$-th measurement with $k=1,\ldots,M_{t}$, which is generated by some $\mathbf{x}_{t,j}$ according to a linear measurement model:
@@
where $\mathbf{v}_{t}$ is zero-mean Gaussian noise with covariance $V$. We denote the concatenation of all measurements by:
Problem.
Given an estimate of the state $\mathbf{x}_{t}$ at time $t$, an input $\mathbf{u}_{t}$, and measurements $\mathbf{z}_{t+1}$, obtain an estimate of $\mathbf{x}_{t+1}$. The association between the measurements $\mathbf{z}_{t+1,k}$ and the variables $\mathbf{x}_{t+1,j}$ that generated them is unknown.
4 Methodology
We formulate a Kalman filter with probabilistic data association to solve the estimation problem defined in Sec. 3.
4.1 Data association
We define the association between a set of measurements $\mathcal{Z}$ and a set of variables $\mathcal{X}$ that generated them as follows.
Definition.
The data association of set $\mathcal{Z}=\{\mathbf{z}_{k}\}_{k=1}^{M}$ to set $\mathcal{X}=\{\mathbf{x}_{j}\}_{j=1}^{N}$ is a function $\delta:\{1,\ldots,M\}\to\{0,1,\ldots,N\}$ that either associates an element $\mathbf{z}_{k}\in\mathcal{Z}$ to an element $\mathbf{x}_{\delta(k)}\in\mathcal{X}$ or indicates via $\delta(k)=0$ that there is no matching element in $\mathcal{X}$.
For example, in MOT, $\mathcal{Z}$ is the set of measurements (e.g., features, bounding boxes, or object segmentations), and $\mathcal{X}$ is the set of tracked objects.
Let $\delta_{t}$ be the data association at time $t$, such that measurement $\mathbf{z}_{t,k}$ is assigned to variable $\mathbf{x}_{t,\delta_{t}(k)}$. We denote the set of all possible data association functions at time $t$ as $\mathcal{D}_{t}=\{\delta_{t}:\{1,...,M_{t}\}\to\{1,...,N\}\}$ and the set of data associations across all times as $\mathcal{D}=\cup_{t=1}^{T}\mathcal{D}_{t}$.
We consider a probabilistic setting, in which we are given a prior probability density $p(\mathbf{x}_{t})$ of the state $\mathbf{x}_{t}$, and aim to compute the posterior density $p(\mathbf{x}_{t+1}|\mathbf{z}_{t+1},\mathbf{u}_{t})$ of $\mathbf{x}_{t+1}$ conditioned on the input $\mathbf{u}_{t}$ and the measurements $\mathbf{z}_{t+1}$. At each time $t$, we treat the data association $\delta_{t}$ as a random variable with a uniform prior $p(\delta_{t})$ independent of $\mathbf{x}_{t}$. Using Bayes rule and assuming the measurements are mutually independent conditioned on the variables that generated them, the density of $\delta_{t}$ conditioned on $\mathbf{z}_{t}$ satisfies:
$p(\delta_{t}|\mathbf{z}_{t})$ $=\frac{p(\mathbf{z}_{t}|\delta_{t})p(\delta_{t})}{p(\mathbf{z}_{t})}\propto p(\mathbf{z}_{t}|\delta_{t})$
[…]
$=\prod_{k=1}^{M_{t}}p\left(\mathbf{z}_{t,k}|\delta_{t}(k)\right).$ (5)
Let $p_{\mathcal{N}}(\cdot;\boldsymbol{\mu},\Sigma)$ denote the density of a Gaussian distribution with mean $\boldsymbol{\mu}$ and covariance $\Sigma$. Since $p(\mathbf{z}_{t,k}|\mathbf{x}_{t,\delta_{t}(k)})=p_{\mathcal{N}}(\mathbf{z}_{t,k};H\mathbf{x}_{t,\delta_{t}(k)},V)$ and $p(\mathbf{x}_{t,\delta_{t}(k)})\allowbreak=p_{\mathcal{N}}(\mathbf{x}_{t,\delta_{t}(k)};\boldsymbol{\mu}_{t,\delta_{t}(k)},\Sigma_{t,\delta_{t}(k)})$, we have:
@@ 4.1 Data association
Next, consider the likelihood that $\mathbf{z}_{t,k}$ is generated by $\mathbf{x}_{t,j}$, which we denote by $w^{t}_{k,j}$. Using (5), we have:
$w^{t}_{k,j}=\!\!\!\!\sum_{\delta\in\mathcal{D}_{t}(k,j)}\!\!p(\delta|\mathbf{z}_{t})=\!\!\!\!\sum_{\delta\in\mathcal{D}_{t}(k,j)}\prod_{k^{\prime}=1}^{M_{t}}p\left(\mathbf{z}_{t,k^{\prime}}|\delta_{t}(k^{\prime})\right),$ (6)
where $\mathcal{D}_{t}(k,j)$ is the set of data association functions that assign measurement $\mathbf{z}_{t,k}$ to variable $\mathbf{x}_{t,j}$. We show that $w^{t}_{k,j}$ can be computed using the permanent of a matrix $Q$ containing the measurement likelihood $p(\mathbf{z}_{t,k}|\mathbf{x}_{t,j})$ as its entries $Q(k,j)$. The permanent of a matrix $Q=\left[Q(k,j)\right]\in\mathbb{R}^{M\times N}$ with $M\leq N$ is defined as:
@@
Let $Q^{t}\in\mathbb{R}^{M_{t}\times N}$ be a matrix with elements $Q^{t}(k,j)=p(\mathbf{z}_{t,k}|\delta_{t}(k)=j)=p(\mathbf{z}_{t,k}|\mathbf{x}_{t,j})$. The data association weight $w^{t}_{k,j}$ in (6) can be computed as:
$w^{t}_{k,j}\propto Q^{t}(k,j)\operatorname*{per}\left(Q^{t}_{-kj}\right),$ (8)
@@
See Supplementary Material Section 7. ∎
Proposition 1 allows us to compute data association weights. Next, we derive a Kalman filter with probabilistic data association.
4.2 Probabilistic data association Kalman filter
@@
Given the prior density $p(\mathbf{x}_{t})$, the input $\mathbf{u}_{t}$, and the measurement $\mathbf{z}_{t+1}$, we aim to find an estimate $q(\bar{\mathbf{x}})$ of the joint posterior $p(\bar{\mathbf{x}}|\mathbf{z}_{t+1},\mathbf{u}_{t})$. To determine $q(\bar{\mathbf{x}})$, we use variational inference [7, Ch. 10], which maximizes the ELBO of the evidence $\log p(\mathbf{z}_{t+1},\mathbf{u}_{t})$ with respect to $q(\bar{\mathbf{x}})$. To save space, we denote $q(\bar{\mathbf{x}})$ by $q$. The evidence can be decomposed as [7, Ch. 10]:
@@
where $\operatorname*{KL}(q||p)$ is the Kullback–Leibler divergence and $\mathcal{L}(q)$ is the ELBO. Since the evidence is constant for given $\mathbf{z}_{t+1},\mathbf{u}_{t}$, the ELBO is maximized when the $\operatorname*{KL}$ term is zero.
To account for probabilistic data association, we formulate the optimization inspired by the EM algorithm [7, Ch. 9]. EM determines the maximum likelihood (ML) or maximum a posterior (MAP) in the presence of unobserved variables. It contains an E-step that computes the expectation w.r.t. the unobserved variables and an M-step that maximizes the log-likelihood. Letting the data association $\delta_{t+1}$ be the unobserved variable, we consider the following optimization problem:
@@
$f^{(i)}(q)=\mathbb{E}_{\delta_{t+1}}\mathbb{E}_{\bar{\mathbf{x}}^{(i)}}\!\left[\mathcal{L}(q,\delta_{t+1})\mid\bar{\mathbf{x}}^{(i)},\mathbf{z}_{t+1}\right],$
$\quad=\mathbb{E}_{\delta_{t+1}}\mathbb{E}_{\bar{\mathbf{x}}^{(i)}}\!\left[\mathbb{E}_{q}\bigl{[}\log\!\frac{p(\bar{\mathbf{x}},\mathbf{z}_{t+1},\mathbf{u}_{t},\delta_{t+1})}{q(\bar{\mathbf{x}})}\bigr{]}\mid\bar{\mathbf{x}}^{(i)}\!,\mathbf{z}_{t+1}\right],$
@@
See Supplementary Material Section 8. ∎
4.3 Comparison with JPDAF
We compare our PKF to the JPDAF [2]. We first show the update scheme difference, then show that the association weights in [2] can be computed using Proposition 1 and compare the tracking errors and speed on simulated data.
4.3.1 Update scheme
Both JPDAF [2] and PKF can be decomposed into two stages. The JPDAF first computes a posterior mean as the normal Kalman filter with each associated measurement, and then averages the posterior means with the association weights and computes the covariance correspondingly. Instead, our PKF first constructs an expanded measurement model and then performs a single update in the same form as the normal Kalman filter but with an expanded measurement vector. The JPDAF maximizes the ELBO with regard to each associated measurement and then averages those estimates, while our PKF maximizes the total weight-averaged ELBO, which does not necessarily maximize each measurement’s individual ELBO but may lead to a higher total ELBO. This difference is underscored by the lower tracking errors in the simulations in Sec. 4.3.3.
Both algorithms can be performed in a coupling or decoupling way. In the coupling way, all the objects are stored in one filter and updated jointly. In the decoupling way, each object is stored and updated separately. Proposition 2 shows the coupling version. In the experiments, we do decoupling updates just like JPDAF. Given measurements and the association weights, the update is the same as Proposition 2 except that there is an estimate of only one object in the state. Regarding the complexity, the Kalman gain in JPDAF is the same as the normal Kalman filter but JPDAF needs all the measurement residuals to update the mean and covariance. In PKF, the complexity of the Kalman gain computation increases due to the expanded measurement model. The original Kalman filter has complexity $\mathcal{O}(m^{2.4}+n^{2})$ [28] in terms of the measurement dimension $m$ and state dimension $n$. Given the number of associated measurements $M_{t}$, the complexity of PKF is $\mathcal{O}\left((M_{t}m)^{2.4}+M_{t}n^{2}\right)$. In practice, $M_{t}$ is usually small because of outlier pruning like validation gating [2] and our ambiguity check in Sec. 5.1. Hence, PKF remains efficient as validated by the experiments in Sec. 4.3.3 and Sec. 5.2.
4.3.2 Weight computation
Following JPDAF [2], we make the assumption that 1) the number of established objects is known and 2) there can be missing or false measurements (clutter) for each object at each time step. Using our notation, the probability of an association event $\delta_{t}$ in [2, Eq. 47] can be written as

### v2

@@ Abstract
In this paper, we derive a new Kalman filter (KF) with probabilistic data association between measurements and states. We formulate a variational inference problem to approximate the posterior density of the state conditioned on the measurement data. We view the unknown data association as a latent variable and apply Expectation Maximization (EM) to obtain a filter with the update step in the same form as the Kalman filter but with an expanded measurement vector of all potential associations. We show that the association probabilities can be computed as permanents of matrices with measurement likelihood entries. We name our probabilistic data association Kalman filter the PKF with P emphasizing both the probabilistic nature of the data association and the matrix permanent computation of the association weights. We compare PKF with the well-established Probabilistic Multi-Hypothesis Tracking (PMHT) and Joint Probabilistic Data Association Filter (JPDAF) in both theory and simulated experiments. The experiments show that we can achieve lower tracking errors than both. We also demonstrate the effectiveness of our filter in multi-object tracking (MOT) on multiple real-world datasets, including MOT17, MOT20, and DanceTrack. We can achieve comparable tracking results with previous KF-based methods without using velocities or doing multi-stage data association and remain real-time. We further show that our PKF can serve as a backbone for other KF-based trackers by applying it to a method that uses varieties of features for association, and improving its results.
Index Terms:
Visual Tracking, Probability and Statistical Methods, Computer Vision for Automation
I Introduction
@@
In estimation tasks where ambiguity exists between measurements and variables of interest, probabilistic data association can prevent catastrophic estimation failures. For example, in multi-object tracking (MOT), there can be a lot of occlusions, causing high ambiguity in data association. An illustration of how probabilistic data association improves MOT is shown in Figure 1. Methods utilizing the Kalman filter (KF) [1, 2, 3, 4] have achieved outstanding performance in MOT but have not considered the impact of probabilistic data association on the tracking process.
@@
We derive a new form of Kalman filter with probabilistic data association. Previous works [5, 6] show that Kalman filter can be derived using variational inference (VI) by maximizing the evidence lower bound (ELBO) of the input and measurement likelihood. To deal with ambiguous associations, we approach the VI problem using Expectation-Maximization (EM) with the data association as the latent variable. In the E-step, we show that association weights can be computed as permanents of matrices with measurement likelihood entries using accelerated algorithms [7, 8]. We also show that the weight computation can be extended to cases with missing or false detections. In the M-step, optimizing the EM objective leads to a Kalman filter with the usual prediction and update steps but with an extended measurement vector and noise covariance matrix capturing all possible data associations. Our formulation is related to Probabilistic Multi-Hypothesis Tracking (PMHT) [9] and the Joint Probabilistic Data Association Filter (JPDAF) [10]. PMHT assumes independence of measurements and tracks in both E-step and M-step, which can easily be violated in some problems. The JPDAF computes a posterior given each associated measurement, i.e., maximizes the ELBO of each measurement, and then computes the weighted average of the posteriors. Our filter, instead, directly optimizes the overall ELBO weighted by the data association weights. We show that our algorithm achieves lower tracking errors than both PMHT and JPDAF, while running at comparable speeds to JPDAF.
Figure 1: Example scene with high ambiguity in DanceTrack [11] with green box detections and orange box tracks. Applying our probabilistic data association Kalman filter on this sequence significantly increases the association quality according to IDF1 [12], and the combined HOTA metric [13].
We also apply our filter to MOT and test our algorithm on multiple real-world datasets, including MOT17 [14], MOT20 [15], and DanceTrack [11]. Without using techniques like multi-stage data association, KF re-updates, object recovery, or considering velocities during association, our method achieves comparable performance to previous KF-based methods [2, 3] while maintaining almost the same inference speed. We further show the compatibility of our PKF to varieties of features for data association by applying it to Hybrid-SORT [4], which uses the techniques from [2, 3] in addition to using detection scores and neural features. By replacing the KF in [4] with our PKF, we can further improve its results. Our contributions are summarized as follows.
@@
  • We show that association probabilities can be computed using matrix permanents, which can accelerate the association weight computation.
  • We demonstrate the effectiveness of the PKF in comparison to PMHT and JPDAF on simulated data. On multiple real-world MOT datasets, PKF achieves comparable results with previous state-of-the-art algorithms by associating only bounding boxes, and can improve the results by replacing the KF part, and runs at 250+ fps on a single laptop CPU, given offline detections [^1] Code is at https://github.com/hwcao17/pkf..
II Related work
II-A Probabilistic data association
@@
II-B Multi-object tracking
State-of-the-art MOT algorithms are dominated by the tracking-by-detection approach, which consists of two steps: object detection [29, 30] and data association. SORT [1] uses the Hungarian algorithm [31] to associate bounding-box object detections and updates the object tracks using the Kalman filter to achieve real-time online tracking. DeepSORT [32] extends SORT [1] by introducing deep visual features for data association. ByteTrack [2] makes an observation that the detectors can make imperfect predictions in complex scenes and proposes a second round of data association for low-confidence detections. OC-SORT [3] analyzes the limitations of SORT [3], namely sensitivity to state noise and temporal error magnification, and proposes an observation-centric re-update step and object momentum for data association.
@@ II-B Multi-object tracking
III Problem formulation
Consider $N$ time-varying variables that need to be estimated, denoted as $\mathbf{x}_{t,j}\in\mathbb{R}^{n}$, $j=1,\ldots,N$. We refer to their concatenation into a single vector at time $t$:
@@ III Problem formulation
where $\mathbf{u}_{t}\in\mathbb{R}^{l}$ is the known input and $\mathbf{w}_{t}$ is zero-mean Gaussian noise with covariance $W$. At each time $t$, we receive $M_{t}$ measurements
The data association between $\mathbf{z}_{t}$ and $\mathbf{x}_{t}$ is defined as follows.
Definition 1.
The data association of $\mathbf{z}_{t}$ to $\mathbf{x}_{t}$ is a function $\delta_{t}:\{1,\ldots,M_{t}\}\to\{0,1,\ldots,N\}$ that either associates an element $\mathbf{z}_{t,k}$ to an element $\mathbf{x}_{t,\delta_{t}(k)}$ or indicates via $\delta_{t}(k)=0$ that there is no matching element in $\mathbf{x}_{t}$.
Given association $\delta_{t}(k)=j$, the measurement model is
@@
where $\mathbf{v}_{t}$ is zero-mean Gaussian noise with covariance $V$.
Problem 1.
Given an estimate of the state $\mathbf{x}_{t}$ at time $t$, an input $\mathbf{u}_{t}$, and measurements $\mathbf{z}_{t+1}$, obtain an estimate of $\mathbf{x}_{t+1}$ with unknown association between the measurements $\mathbf{z}_{t+1,k}$ and the variables $\mathbf{x}_{t+1,j}$.
IV Methodology
We formulate a Kalman filter with probabilistic data association to solve the estimation problem defined in Sec. III.
IV-A Data association
Following Definition 1, $\delta_{t}$ is the data association at time $t$ indicating that measurement $\mathbf{z}_{t,k}$ is assigned to variable $\mathbf{x}_{t,\delta_{t}(k)}$. We denote the set of all possible association functions at time $t$ as $\mathcal{D}_{t}=\{\delta_{t}:\{1,...,M_{t}\}\to\{1,...,N\}\}$ and the set of data associations across all times as $\mathcal{D}=\cup_{t=1}^{T}\mathcal{D}_{t}$.
We consider a probabilistic setting, in which we are given a prior probability density $p(\mathbf{x}_{t})$ of the state $\mathbf{x}_{t}$, and aim to compute the posterior density $p(\mathbf{x}_{t+1}|\mathbf{z}_{t+1},\mathbf{u}_{t})$ of $\mathbf{x}_{t+1}$ conditioned on the input $\mathbf{u}_{t}$ and the measurements $\mathbf{z}_{t+1}$. At each time $t$, we treat the data association $\delta_{t}$ as a random variable with a uniform prior $p(\delta_{t})$ independent of $\mathbf{x}_{t}$. Using Bayes’ rule and assuming the measurements are mutually independent conditioned on the variables that generated them, the density of $\delta_{t}$ conditioned on $\mathbf{z}_{t}$ satisfies:
$p(\delta_{t}|\mathbf{z}_{t})\!=\!\frac{p(\mathbf{z}_{t}|\delta_{t})p(\delta_{t})}{p(\mathbf{z}_{t})}\!\propto\!p(\mathbf{z}_{t}|\delta_{t})\!=\!\prod_{k=1}^{M_{t}}\!p\left(\mathbf{z}_{t,k}|\delta_{t}(k)\right).$ (5)
Let $p_{\mathcal{N}}(\cdot;\boldsymbol{\mu},\Sigma)$ denote the density of a Gaussian distribution with mean $\boldsymbol{\mu}$ and covariance $\Sigma$. Assuming Gaussian prior $p(\mathbf{x}_{t,\delta_{t}(k)})\allowbreak=p_{\mathcal{N}}(\mathbf{x}_{t,\delta_{t}(k)};\boldsymbol{\mu}_{t,\delta_{t}(k)},\Sigma_{t,\delta_{t}(k)})$ and measurement $p(\mathbf{z}_{t,k}|\mathbf{x}_{t,\delta_{t}(k)})=p_{\mathcal{N}}(\mathbf{z}_{t,k};H\mathbf{x}_{t,\delta_{t}(k)},V)$, we have:
@@ IV-A Data association
Consider the likelihood that a specific measurement $\mathbf{z}_{t,k}$ is generated by $\mathbf{x}_{t,j}$, denoted by $w^{t}_{k,j}$. Using (5), we have:
$w^{t}_{k,j}=\!\!\!\!\sum_{\delta_{t}\in\mathcal{D}_{t}(k,j)}\!\!p(\delta_{t}|\mathbf{z}_{t})=\!\!\!\!\sum_{\delta_{t}\in\mathcal{D}_{t}(k,j)}\prod_{r=1}^{M_{t}}p\left(\mathbf{z}_{t,r}|\delta_{t}(r)\right),$ (6)
where $\mathcal{D}_{t}(k,j)$ is the set of data association functions that assign measurement $\mathbf{z}_{t,k}$ to variable $\mathbf{x}_{t,j}$. We show that $w^{t}_{k,j}$ can be computed using the permanent of a matrix $Q$ containing the measurement likelihood $p(\mathbf{z}_{t,k}|\mathbf{x}_{t,j})$ as its entries $Q(k,j)$.
Definition 2.
The permanent of a matrix $Q=\left[Q(k,j)\right]\in\mathbb{R}^{M\times N}$ with $M\leq N$ is:
@@
Given this definition, we have the following proposition to compute the data association weights.
@@
Let $Q^{t}\in\mathbb{R}^{M_{t}\times N}$ be a matrix with elements $Q^{t}(k,j)=p(\mathbf{z}_{t,k}|\delta_{t}(k)=j)=p(\mathbf{z}_{t,k}|\mathbf{x}_{t,j})$. The data association weight $w^{t}_{k,j}$ in can be computed as:
$w^{t}_{k,j}\propto Q^{t}(k,j)\operatorname*{per}\left(Q^{t}_{-kj}\right),$
@@
See Supplementary Material Section VII. ∎
Proposition 1 allows us to compute data association weights. Next, we derive our PKF.
IV-B Probabilistic data association Kalman filter
@@
Given the prior density $p(\mathbf{x}_{t})$, the input $\mathbf{u}_{t}$, and the measurement $\mathbf{z}_{t+1}$, we aim to find an estimate $q(\bar{\mathbf{x}})$ of the joint posterior $p(\bar{\mathbf{x}}|\mathbf{z}_{t+1},\mathbf{u}_{t})$. To determine $q(\bar{\mathbf{x}})$, we use variational inference [40, Ch. 10], which maximizes the lower bound of the evidence $\log p(\mathbf{z}_{t+1},\mathbf{u}_{t})$ with respect to $q(\bar{\mathbf{x}})$. To save space, we denote $q(\bar{\mathbf{x}})$ by $q$. The evidence can be decomposed as [40, Ch. 10]:
@@
where $\mathcal{L}(q)$ is the ELBO (evidence lower bound) since the Kullback–Leibler divergence $\operatorname*{KL}(q||p)\geq 0$. Following standard practice [40, Ch. 10], we optimize the ELBO
$q^{*}\!\in\!\arg\max_{q}\mathcal{L}(q)\!=\!\arg\max_{q}\mathbb{E}_{q}\bigl{[}\log\frac{p(\bar{\mathbf{x}},\mathbf{z}_{t+1},\mathbf{u}_{t})}{q(\bar{\mathbf{x}})}\bigr{]}.$ (11)
To account for probabilistic data association, we formulate the optimization by the EM algorithm [40, Ch. 9]. EM determines the maximum likelihood (ML) or maximum a posteriori (MAP) in the presence of unobserved variables. It contains an E-step that computes the expectation w.r.t. the unobserved variables and an M-step that maximizes the log-likelihood. Letting the data association $\delta_{t+1}$ be the unobserved variable, we consider the following problem:
@@
$=\arg\max_{q}\mathbb{E}_{\delta_{t+1}}\mathbb{E}_{\bar{\mathbf{x}}^{(i)}}\!\left[\mathcal{L}(q,\delta_{t+1})\mid\bar{\mathbf{x}}^{(i)},\mathbf{z}_{t+1}\right]$
$=\arg\max_{q}\mathbb{E}_{\delta_{t+1}}\mathbb{E}_{\bar{\mathbf{x}}^{(i)}}\!\!\left[\mathbb{E}_{q}\bigl{[}\log\!\frac{p(\bar{\mathbf{x}},\mathbf{z}_{t+1},\mathbf{u}_{t},\delta_{t+1})}{q(\bar{\mathbf{x}})}\bigr{]}\!\!\mid\!\!\bar{\mathbf{x}}^{(i)}\!,\mathbf{z}_{t+1}\!\right]$
$=\arg\max_{q}\sum_{k=1}^{M_{t}}\sum_{j=1}^{N}\omega^{t+1}_{kj}\log p(\mathbf{z}_{t+1,k}|\mathbf{x}_{t+1,j}),$
@@
See Supplementary Material Section VIII. ∎
IV-C Comparison with PMHT and JPDAF
We discuss the relationship of PKF to PMHT [9] and JPDAF [10] and compare them in simulated experiments.
IV-C1 Comparison with PMHT
The difference between our PKF and PMHT [9] is that PMHT assumes independence in both the E-step (association) and M-step (update), while the PKF does not. During association, PMHT computes the probability of each measurement and object individually without considering if the measurement is already assigned to another object. In contrast, PKF does the association in a joint manner by evaluating probabilities of the injective association functions, where a measurement can not be assigned to another object after being assigned to one. During the update step, PMHT updates each object individually, while PKF updates objects jointly, i.e., there are cross-correlation blocks is the object covariances.
IV-C2 Comparison with JPDAF
The major difference is in the update step, we first show the update scheme difference, then show that the association weights in JPDAF can be computed using Proposition 1.
Update scheme Both JPDAF [10] and PKF can be decomposed into two stages. The JPDAF first computes a posterior mean as the normal Kalman filter with each associated measurement, and then averages the posterior means with the association weights and computes the covariance correspondingly. Instead, our PKF first constructs an expanded measurement model and then performs a single update in the same form as the normal Kalman filter but with an expanded measurement vector. The JPDAF maximizes the ELBO with regard to each associated measurement and then averages those estimates, while our PKF maximizes the total weight-averaged ELBO, which does not necessarily maximize each measurement’s individual ELBO but may lead to a higher total ELBO. This difference is underscored by the lower tracking errors in the simulations in Sec. IV-C3.
Both algorithms can be performed in a coupling or decoupling way. In the coupling way, all the objects are stored in one filter and updated jointly. In the decoupling way, each object is stored and updated separately, i.e., setting the cross-correlation blocks of $\Sigma_{t}$ in Proposition 2 to 0. In the experiments, we do decoupling updates just like JPDAF. Regarding the complexity, the Kalman gain in JPDAF is the same as the normal Kalman filter but JPDAF needs all the measurement residuals to update the mean and covariance. In PKF, the complexity of the Kalman gain computation increases due to the expanded measurement model. The original Kalman filter has complexity $\mathcal{O}(m^{2.4}+n^{2})$ [41] in terms of the measurement dimension $m$ and state dimension $n$. Given the number of associated measurements $M_{t}$, the complexity of PKF is $\mathcal{O}\left((M_{t}m)^{2.4}+M_{t}n^{2}\right)$.
Weight computation Following JPDAF [10], we make the assumption that 1) the number of established objects is known and 2) there can be missing or false measurements (clutter). Using our notation, the probability of an association event $\delta_{t}$ in [10, Eq. 47] can be written as